OpenStudy (anonymous):

Determine the length of the leg of a 45 – 45 – 90 triangle with a hypotenuse length of 15 inches.

OpenStudy (anonymous):

i got \[7.5\sqrt{2}\]??

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

hmm....

OpenStudy (lgbasallote):

let's check \[(7.5 \sqrt 2)^2 + (7.5 \sqrt 2)^2 = 15^2\] do you think those two equal?

OpenStudy (anonymous):

yeah?

OpenStudy (lgbasallote):

then you're right